OpenStudy (sboodhoo):

PLEASE HELP WILL FAN AND MEDAL! If y = 5x − 3, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(2, 5), (−3, −12), (−4, −18)} {(−1, −8), (0, −3), (3, 12)} {(1, 4), (7, 25), (3, 10)} {(5, 20), (6, 12), (1, 9)}

OpenStudy (seanstewart16):

you want me to explain how to do it, or do it for you?

OpenStudy (sboodhoo):

both kinda?

OpenStudy (seanstewart16):

alright.

OpenStudy (seanstewart16):

this might take a few minutes

OpenStudy (sboodhoo):

k

OpenStudy (seanstewart16):

take the original equation y=5x-3, and you need to plug all of those sets in to see if they apply to the origin

OpenStudy (seanstewart16):

the first set, plug in 2 for x and 5 for y, and let me know what yop got

OpenStudy (sboodhoo):

10, 15????????

OpenStudy (seanstewart16):

no, plug them in like this. y=5x-3 so that 5=5(2)-3

OpenStudy (seanstewart16):

plug them in like youre checking points on a graph

OpenStudy (sboodhoo):

?

OpenStudy (seanstewart16):

just o you know, i workrd it out, its the second one. but you still need to understand why

OpenStudy (seanstewart16):

ok, look. ill just show you why it is the second one, ok?

OpenStudy (seanstewart16):

the first set is (-1,-8), so we must plug it in to see if it works. plug in -1 for x and -8 for y -8=5(-1)-3